# A Single-Step Non-Autoregressive Automatic Speech Recognition Architecture with High Accuracy and Inference Speed

### v1

@@ Abstract
Non-autoregressive (NAR) automatic speech recognition (ASR) models predict tokens independently and simultaneously, bringing high inference speed. However, there is still a gap in the accuracy of the NAR models compared to the autoregressive (AR) models. To further narrow the gap between the NAR and AR models, we propose a single-step NAR ASR architecture with high accuracy and inference speed, called EfficientASR. It uses an Index Mapping Vector (IMV) based alignment generator to generate alignments during training, and an alignment predictor to learn the alignments for inference. It can be trained end-to-end (E2E) with cross-entropy loss combined with alignment loss. The proposed EfficientASR achieves competitive results on the AISHELL-1 and AISHELL-2 benchmarks compared to the state-of-the-art (SOTA) models. Specifically, it achieves character error rates (CER) of 4.26%/4.62% on the AISHELL-1 dev/test dataset, which outperforms the SOTA AR Conformer with about 30x inference speedup.
keywords:
ASR, E2E, non-autoregressive, single-step NAR, EfficientASR
1 Introduction
Over the past few years, end-to-end (E2E) models have achieved remarkable results on ASR tasks. There are three popular E2E approaches: connectionist temporal classification (CTC) [1], recurrent neural network transducer (RNN-T) [2, 3], and attention-based encoder-decoder (AED) [4, 5]. Among them, the CTC model is non-autoregressive. Although NAR models do not perform as well as AR models due to the conditional independence assumption in decoding, their parallel decoding can greatly reduce the decoding time. To bridge the gap between NAR and AR models, much work has been done.
@@ 1 Introduction
Depending on the number of iterations in inference, NAR models are classified as iterative or single-step models. Among the former, A-FMLM [6], a non-autoregressive Transformer model (NAT), is the first attempt to introduce the conditional masked language model (CMLM) [7] into ASR. A-FMLM is designed to predict masked tokens conditioned on unmasked ones and whole speech embeddings. However, its target token length needs to be pre-defined, which limits the performance of the model. To address this issue, ST-NAT [8] uses a CTC module to predict the length of the target sequence. Unlike ST-NAT, Mask-CTC and its variants [9, 10, 11] propose to use the CMLM decoder to refine CTC decodings.
@@
However, iterative models require multiple iterations to achieve a competitive result, limiting the speed of inference in practice. To overcome this limitation, single-step NAR models are proposed. LASO [12] implements parallel decoding of Transformer-style models based on the position-dependent summarizer (PDS) module, but it requires a predefined token length. InterCTC [13, 14] uses an additional intermediate loss of CTC to relax the conditional independence assumption of CTC models, improving the accuracy while maintaining the inference speed of the CTC model. CASS-NAT and its variants [15, 16, 17, 18] conduct in-depth exploration on the combination of CTC alignment and NAT, further improving the performance of the single-step NAR models. Different from the CTC-based or NAT-based NAR models mentioned above, Continuous Integrate-and-Fire (CIF) [19] is the first attempt to model ASR tasks by explicitly predicting the length of the target sequence. However, due to the conditional independence assumption, the accuracy of these mentioned NAR models is significantly inferior to the SOTA AR models.
@@
Paraformer [20], a CIF-based model, designs a glancing language model (GLM) based sampler to strengthen the NAR decoder with the ability to model token interdependency. According to the Paraformer’s report, it can achieve comparable performance to that of the SOTA AR transformer on open source Mandarin Chinese datasets [20]. Despite the success of Paraformer in NAR ASR modeling, it requires a well-designed two-pass training process, which makes the model less compact and less efficient.
@@
Inspired by EfficientTTS [21], a NAR Text-to-Speech (TTS) model, we propose a single-step NAR ASR architecture, termed EfficientASR. The EfficientASR is more compact compared to other high-performance NAR models, with efficient training and inference processes. And it achieves fairly competitive results on the public AISHELL-1 and AISHELL-2 benchmarks compared to the SOTA models.
2 Methods
The main contribution of this paper is to introduce the methods from EfficientTTS [21] into NAR ASR modeling. For the ASR task, we redesigned the model structure of EfficientTTS [21] and proposed EfficientASR. In this section, we will illustrate the structure of the proposed EfficientASR.
@@ 2 Methods
Figure 1: The model architecture of the proposed EfficientASR. ($\bigotimes$ represents the matrix multiplication operation)
2.1 Overview
@@
The architecture of the proposed EfficientASR is shown in Fig. 1. The training network consists of six modules, namely mel-encoder, text-encoder, alignment predictor, alignment generator, attention reconstruction, and decoder. The text-encoder and alignment generator are designed for training and are no longer activated during inference.
@@
Let the acoustic feature sequence be $\bm{x}\in\mathcal{R}^{T}$ and the transcription sequence be $\bm{y}\in\mathcal{R}^{L}$, where $T$ and $L$ are the length of acoustic feature sequence and transcription sequence respectively. During training, the mel-encoder encodes $\bm{x}$ into the acoustic embeddings $\bm{e}^{(\mathrm{s})}\in\mathcal{R}^{T,d}$, and the text-encoder encodes $\bm{y}$ into text embeddings $\bm{e}^{(\mathrm{t})}\in\mathcal{R}^{L,d}$, where $d$ is the embedding dimension. The alignment generator leverages $\bm{e}^{(\mathrm{s})}$ and $\bm{e}^{(\mathrm{t})}$ to generate the alignment $\bm{\delta}\in\mathcal{R}^{T}$ between acoustic and text embeddings. $\bm{\delta}$ is then passed to the attention reconstruction module to construct an attention matrix $\bm{\hat{\alpha}}\in\mathcal{R}^{L,T}$. Multiplying by $\bm{\hat{\alpha}}$, the acoustic embeddings $\bm{e}^{(\mathrm{s})}$ are transformed into semantic encodings of length equal to the output tokens. The semantic encodings are finally fed into a decoder to predict tokens $\bm{y}^{*}\in\mathcal{R}^{L}$. However, at the inference stage, the model cannot obtain alignments through the generator because of the lack of transcriptions. Therefore, during training, we use the predictor to learn the alignment produced by the generator. Then during inference, the output $\bm{\delta^{*}}$ of the alignment predictor is used to reconstruct the attention matrix and perform the inference of the model.
@@
The mel-encoder is the same as the AR Conformer encoder [22], consisting of several conformer blocks. The text-encoder and the decoder are built with Transformer [23] encoder blocks. Other modules will be detailed in the following parts.
@@
2.2 Alignment generator
The alignment generator is built based on the Index Mapping Vector (IMV), proposed in [21], which provides a creative way to describe the alignment between input and output sequences. The calculation process of IMV-based alignment for ASR is depicted in Fig. 2(a). First, a scaled dot-product attention matrix $\bm{\alpha}\in\mathcal{R}^{T,L}$ is built by Eq. (1), with $\bm{e}^{(\mathrm{s})}$ as Query and $\bm{e}^{(\mathrm{t})}$ as Key:
$\alpha_{i,j}=\frac{\mathrm{exp}((\bm{e}_{i}^{(\mathrm{s})}\cdot\bm{e}_{j}^{(\mathrm{t})})*d^{-0.5})}{\sum_{j=0}^{L-1}\mathrm{exp}((\bm{e}_{i}^{(\mathrm{s})}\cdot\bm{e}_{j}^{(\mathrm{t})})*d^{-0.5})}$ (1)
where $10\leq i\leq T-$ 1, $10\leq j\leq L-$ 1, and $d$ is the dimension of $\bm{e}^{(\mathrm{s})}$ and $\bm{e}^{(\mathrm{t})}$. Then the IMV $\bm{p}$ is calculated by Eq. (2):
$\bm{p}=\bm{\alpha}\cdot\bm{I}$ (2)
where $\bm{I}=[0,1,...,L-1]$ is the index vector of transcription, and $\Sigma_{j=0}^{L-1}\alpha_{i,j}=1$. $p_{i}=\Sigma_{j=0}^{L-1}(\alpha_{i,j}*j)$, which is the weighted sum of the attention weight between the $i$-th input frame and each output token and the position index of each output token. $p_{i}$ can be seen as the expected alignment location of the $i$-th input frame in the length range of the output text. We define the increment of alignment locations of adjacent audio frames in the transcription as ${\Delta p_{i}}$, which be calculated by Eq. (3). According to the monotonicity assumption in EfficientTTS [21], when there is a monotonic correspondence between audio and transcription, it can be deduced that $\Delta p_{i}\geq 0$, and the proof process can be found in [21].
$\Delta p_{i}=p_{i}-p_{i-1},1\leq i\leq T-1$ (3)
However, at the beginning of training, the model has not been trained well enough to satisfy $\Delta p_{i}\geq 0$ automatically. To ensure this monotonic relationship, $\Delta p_{i}$ is activated by ReLU, as shown in Eq. (4):
$\delta_{i}=\left\{\begin{array}[]{ll}0,i=0\\ {\rm ReLU}(\Delta p_{i}),1\leq i\leq T-1\end{array}\right.$ (4)
where $\delta_{0}$ is set to 00. $\delta_{i}$ can be understood as the weight of each encoder step involved in the calculation of the decoding label.
Table 1: CER(%) and RTF results of our models on AISHELL-1 and AISHELL-2 tasks and comparison with previous work. (†: RTF is evaluated with batchsize of 8, ‡: Model parameters in the inference phase.)
@@ 2.2 Alignment generator
EfficientASR Base (Ours) w/o 4.30 4.66 6.03 43.6M ‡ 0.0058
EfficientASR Large (Ours) w/o 4.26 4.62 5.76 76.0M ‡ 0.0066
2.3 Attention reconstruction
After obtaining alignment $\bm{\delta}$, our goal is to reconstruct an attention matrix, such that the acoustic embeddings $\bm{e}^{(\mathrm{s})}$ of length $T$ can be transformed into semantic encodings of length $L$ by multiplying with the attention matrix. We follow the similar idea of [21] in reconstructing the attention matrix and the construction process of the attention matrix can be summarized as shown in Fig. 2(b). First, a monotonically increasing alignment position vector $\bm{p\mathrm{{}^{\prime}}}$ is generated by accumulating $\delta_{i}$, as shown in Eq. (5). Then, a scaling strategy as shown in Eq. (6) is applied to ensure that the maximum value of $\bm{p^{\prime}}$ is equal to the length of the output tokens (since the token index starts from 0, in practice, take $L-1$).
$p_{i}{{}^{\prime}}=\sum_{m=0}^{i}\delta_{m},0\leq i\leq T-1$ (5)
$\hat{p}_{i}=\frac{p_{i}{{}^{\prime}}-p_{0}{{}^{\prime}}}{p_{T-1}{{}^{\prime}}-p_{0}{{}^{\prime}}}*(L-1),0\leq i\leq T-1$ (6)
Since both $\bm{\hat{p}}$ and transcription index vector $\bm{I}$ are monotonic sequences ranging from 00 to $L-1$, a distance-aware attention reconstruction is performed by Eq. (7):
@@ 2.3 Attention reconstruction
where $d_{i,j}=(\hat{p}_{i}-I_{j})^{2},0\leq i\leq T-1,0\leq j\leq L-1$, represents the semantic distance between the $i$-th encoder frame and the $j$-th token. $\sigma$ is a learnable hyperparameter. Obviously, the smaller distance $d_{i,j}$ is, the larger attention weight $\hat{\alpha}_{i,j}$.
2.4 Alignment predictor
As mentioned above, an alignment predictor is required to generate alignment at inference stage. We built it with two 1D-Convolution layers, each followed by the layer normalization and ReLU function. During training, it takes acoustic embeddings $\bm{e}^{(\mathrm{s})}$ as inputs to generate a predicted alignment $\bm{\delta}^{*}$, while using alignment $\bm{\delta}$ as the label to update the predictor. In the inference stage, we use the alignment $\bm{\delta}^{*}$ produced by the predictor to construct the attention matrix, as shown in the Fig. 1(b). At the same time, we accumulate the $\bm{\delta}^{*}$ and round it to predict the number of output tokens. The Mean Square Error (MSE) loss is used to learn the alignment, and the total loss is defined as:
$\mathcal{L}_{total}=\mathcal{L}_{\rm{CE}}(\bm{y},\bm{y^{*}})+\lambda\mathcal{L}_{\rm{MSE}}(\bm{\delta},\bm{\delta^{*}})\$ (8)
where $\lambda$ is the weight of alignment loss, and $\mathcal{L}_{\rm{CE}}$ is the cross-entropy (CE) loss between the outputs and labels.
3 Experiments
3.1 Experimental Setup
The proposed method is evaluated on two public Mandarin corpus: 178 hours AISHELL-1 [26] and 1000 hours AISHELL-2 [27]. We use 80-channel filter banks computed from a 25 ms window with a stride of 10 ms as features. SpecAugment [28] and speed perturbation are used for data augmentation for all experiments. The sizes of the vocabulary for the AISHELL-1 and AISHELL-2 tasks are $4,233$ and $5,208$, respectively. The EfficientASR model is built in base and large size. For both sizes, the layers of {mel-encoder, text-encoder, alignment predictor, decoder} are set to {12, 1, 2, 6}, while attention head of base model is $4$ with hidden dimension $256$ and attention head of large model is $6$ with hidden dimension $384$. The hyperparameter $\lambda$ in Eq. (8) is set to $1.0$ and the learnable hyperparameter $\sigma$ in Eq. (7) is initialized with $0.5$. The proposed models are developed based on ESPnet [29]  [^1] https://github.com/espnet/espnet.git, and other configurations can be found there. All experiments were performed on $8$ NVIDIA Tesla V100 GPUs.

### v4

# EffectiveASR: A Single-Step Non-Autoregressive Mandarin Speech Recognition Architecture with High Accuracy and Inference Speed

@@ Abstract
Non-autoregressive (NAR) automatic speech recognition (ASR) models predict tokens independently and simultaneously, bringing high inference speed. However, there is still a gap in the accuracy of the NAR models compared to the autoregressive (AR) models. In this paper, we propose a single-step NAR ASR architecture with high accuracy and inference speed, called EffectiveASR. It uses an Index Mapping Vector (IMV) based alignment generator to generate alignments during training, and an alignment predictor to learn the alignments for inference. It can be trained end-to-end (E2E) with cross-entropy loss combined with alignment loss. The proposed EffectiveASR achieves competitive results on the AISHELL-1 and AISHELL-2 Mandarin benchmarks compared to the leading models. Specifically, it achieves character error rates (CER) of 4.26%/4.62% on the AISHELL-1 dev/test dataset, which outperforms the AR Conformer with about 30x inference speedup.
Index Terms:
ASR, single-step NAR, EffectiveASR
I Introduction
In recent years, end-to-end (E2E) models have outperformed traditional hybrid systems in automatic speech recognition (ASR) tasks. Among the three widely used E2E methods—connectionist temporal classification (CTC) [1], recurrent neural network transducer (RNN-T) [2, 3], and attention-based encoder-decoder (AED) [4, 5, 6]—AED models have become the preferred choice for sequence-to-sequence ASR modeling, thanks to their superior recognition accuracy, as seen in models like Transformer and Conformer. However, despite their strong performance, AED models utilize an auto-regressive (AR) decoder that generates tokens sequentially, with each token dependent on the preceding ones. This results in computational inefficiency, as decoding time increases linearly with the length of the output sequence. To address this issue and speed up inference, non-autoregressive (NAR) models have been introduced, allowing for parallel generation of output sequences.
@@ I Introduction
Depending on the number of iterations in inference, NAR models are classified as iterative or single-step models. Among the former, A-FMLM [7], a non-autoregressive Transformer model (NAT), is the first attempt to introduce the conditional masked language model (CMLM) [8] into ASR. A-FMLM is designed to predict masked tokens conditioned on unmasked ones and whole speech embeddings. However, its target token length needs to be pre-defined, which limits the model’s performance. To address this issue, ST-NAT [9] uses a CTC module to predict the length of the target sequence. Unlike ST-NAT, Mask-CTC and its variants [10, 11, 12] propose to use the CMLM decoder to refine CTC.
@@
However, iterative models require multiple iterations to achieve a competitive result, limiting the speed of inference in practice. To overcome this limitation, single-step NAR models are proposed. LASO [13] implements parallel decoding of Transformer-style models based on the position-dependent summarizer (PDS) module, but it requires a predefined token length. InterCTC [14, 15] uses an additional intermediate loss of CTC to relax the conditional independence assumption of CTC models, improving the accuracy while maintaining the inference speed of the CTC model. CASS-NAT and its variants [16, 17, 18, 19] conduct in-depth exploration on the combination of CTC alignment and NAT, further improving the performance of the single-step NAR models. Different from the CTC-based or NAT-based NAR models mentioned above, Continuous Integrate-and-Fire (CIF) [20] is the first attempt to model ASR tasks by explicitly predicting the length of the target sequence. However, due to the conditional independence assumption, the accuracy of these mentioned NAR models is significantly inferior to the leading AR models.
@@
Paraformer [21], a CIF-based model, designs a glancing language model (GLM) based sampler to strengthen the NAR decoder with the ability to model token interdependency. According to the Paraformer’s report, it can achieve comparable performance to that of the AR transformer on open source Mandarin Chinese datasets [22, 23]. A step further, E-Paraformer [24] proposes a novel monotonic alignment mechanism called Parallel Integrate-and-Fire (PIF), which not only enables parallel computation instead of the recursive mechanism in CIF, but also incorporates global context modeling capabilities and gets both better CER performance and faster speeds than the Paraformer. Despite the success of Paraformer and E-Paraformer in NAR ASR modeling, they both require a well-designed two-pass training process, which makes the model less compact and less efficient.
@@
In this paper, we propose a single-step NAR ASR architecture inspired by the Hard Monotonic Alignment (HMA) method applied in [25, 26], termed EffectiveASR. The EffectiveASR is more compact compared to other high-performance NAR models, with efficient training and inference processes. And it achieves fairly competitive results on the public AISHELL-1 and AISHELL-2 benchmarks compared to the leading models.
II Methods
@@
Figure 1: The model architecture of the proposed EffectiveASR. ($\bigotimes$ represents the matrix multiplication operation)
II-A Overview
@@
The architecture of the proposed EffectiveASR is shown in Fig. 1. The training network consists of six modules, namely mel-encoder, text-encoder, alignment predictor, alignment generator, attention reconstruction, and decoder. The text-encoder and alignment generator are designed for training and are no longer activated during inference.
@@
Let the acoustic feature sequence be $\boldsymbol{x}\in\mathcal{R}^{T}$ and the transcription sequence be $\boldsymbol{y}\in\mathcal{R}^{L}$, where $T$ and $L$ are the length of acoustic feature sequence and transcription sequence respectively. During training, the mel-encoder encodes $\boldsymbol{x}$ into the acoustic embeddings $\boldsymbol{e}^{(\mathrm{s})}\in\mathcal{R}^{T,d}$, and the text-encoder encodes $\boldsymbol{y}$ into text embeddings $\boldsymbol{e}^{(\mathrm{t})}\in\mathcal{R}^{L,d}$, where $d$ is the embedding dimension. The alignment generator leverages $\boldsymbol{e}^{(\mathrm{s})}$ and $\boldsymbol{e}^{(\mathrm{t})}$ to generate the alignment $\boldsymbol{\delta}\in\mathcal{R}^{T}$ between acoustic and text embeddings. $\boldsymbol{\delta}$ is then passed to the attention reconstruction module to construct an attention matrix $\boldsymbol{\hat{\alpha}}\in\mathcal{R}^{L,T}$. By multiplying with $\boldsymbol{\hat{\alpha}}$, the acoustic embeddings $\boldsymbol{e}^{(\mathrm{s})}$ are transformed into semantic encodings, with a length corresponding to that of the output tokens. The semantic encodings are finally fed into the decoder to predict tokens $\boldsymbol{y}^{*}\in\mathcal{R}^{L}$. However, at the inference stage, the model cannot obtain alignments through the generator because of the lack of transcriptions. Therefore, during training, we use the predictor to learn the alignment produced by the generator. Then during inference, the output $\boldsymbol{\delta^{*}}$ of the alignment predictor is used to reconstruct the attention matrix and perform the inference of the model.
@@
The mel-encoder is the same as the Conformer encoder [27], consisting of several conformer blocks. The text-encoder and decoder are built with Transformer [28] encoder blocks. Other modules will be detailed in the following parts.
@@
II-B Alignment generator
The alignment generator is built based on the Index Mapping Vector (IMV), proposed in [25], which provides a creative way to describe the alignment between input and output sequences. The calculation process of IMV-based alignment for ASR is depicted in Fig. 2(a). First, a scaled dot-product attention matrix $\boldsymbol{\alpha}\in\mathcal{R}^{T,L}$ is built by Eq. (1), with $\boldsymbol{e}^{(\mathrm{s})}$ as Query and $\boldsymbol{e}^{(\mathrm{t})}$ as Key:
$\alpha_{i,j}=\frac{\mathrm{exp}((\boldsymbol{e}_{i}^{(\mathrm{s})}\cdot\boldsymbol{e}_{j}^{(\mathrm{t})})*d^{-0.5})}{\sum_{j=0}^{L-1}\mathrm{exp}((\boldsymbol{e}_{i}^{(\mathrm{s})}\cdot\boldsymbol{e}_{j}^{(\mathrm{t})})*d^{-0.5})}$ (1)
where $10\leq i\leq T-$ 1, $10\leq j\leq L-$ 1, and $d$ is the dimension of $\boldsymbol{e}^{(\mathrm{s})}$ and $\boldsymbol{e}^{(\mathrm{t})}$. Then the IMV $\boldsymbol{p}$ is calculated by Eq. (2):
$\boldsymbol{p}=\boldsymbol{\alpha}\cdot\boldsymbol{I}$ (2)
where $\boldsymbol{I}=[0,1,...,L-1]$ is the index vector of transcription, and $\Sigma_{j=0}^{L-1}\alpha_{i,j}=1$. Here $p_{i}=\Sigma_{j=0}^{L-1}(\alpha_{i,j}*j)$ is the weighted sum of the attention weight between the $i$-th input frame and each output token and the position index of each output token. $p_{i}$ can be seen as the expected alignment location of the $i$-th input frame in the length range of the output text. We define the increment of alignment locations of adjacent audio frames in the transcription as ${\Delta p_{i}}$, which is calculated by Eq. (3). According to the monotonicity assumption in [25, 26, 6], when there is a monotonic correspondence between audio and transcription, it can be deduced that $\Delta p_{i}\geq 0$, and the proof process can be found in [25, 6].
$\Delta p_{i}=$ $p_{i}-p_{i-1},$ $1\displaystyle 1\leq i\leq T-$ 1 (3)
However, at the beginning of training, the model has not been trained well enough to satisfy $\Delta p_{i}\geq 0$ easily. To guarantee a strictly monotonic constraint, $\Delta p_{i}$ is activated by ReLU, as shown in Eq. (6):
$\delta_{i}=\left\{\begin{array}[]{ll}0,&i=0\\ {\rm ReLU}(\Delta p_{i}),&1\leq i\leq T-1\end{array}\right.$ (6)
where $\delta_{0}$ is set to 00. $\delta_{i}$ can be seen as the weight of each encoder step involved in the calculation of decoding label.
TABLE I: CER(%) and RTF results of our models on AISHELL-1 and AISHELL-2 tasks and the comparison with previous AR and NAR works. (†: RTF is evaluated with batchsize of 8, ‡: Model parameters in the inference phase.)
@@ II-B Alignment generator
E-Paraformer [24] w/o 4.36 4.79 6.08 43.6M 0.0069
EffectiveASR Base (Ours) w/o 4.30 4.66 6.03 43.6M ‡ 0.0058
EffectiveASR Large (Ours) w/o 4.26 4.62 5.76 76.0M ‡ 0.0066
II-C Attention reconstruction
After obtaining alignment $\boldsymbol{\delta}$, our goal is to reconstruct an attention matrix, such that the acoustic embeddings $\boldsymbol{e}^{(\mathrm{s})}$ of length $T$ can be transformed into semantic encodings of length $L$ by multiplying with the attention matrix. We follow the similar idea of [25, 26] in reconstructing the attention matrix and the construction process of the attention matrix can be summarized as shown in Fig. 2(b). First, a monotonically increasing alignment position vector $\boldsymbol{p\mathrm{{}^{\prime}}}$ is generated by accumulating $\delta_{i}$, as shown in Eq. (7). Next, a scaling strategy, as described in Eq. (8), is applied to ensure that the maximum value of $\boldsymbol{p^{\prime}}$ matches the length of the output tokens.
$p_{i}{{}^{\prime}}=$ $\sum_{m=0}^{i}\delta_{m},$ $1\displaystyle 0\leq i\leq T-$ 1 (7)
$\hat{p}_{i}=$ $\frac{p_{i}{{}^{\prime}}-p_{0}{{}^{\prime}}}{p_{T-1}{{}^{\prime}}-p_{0}{{}^{\prime}}}*(L-1),$ $1\displaystyle 0\leq i\leq T-$ 1 (8)
Since both $\boldsymbol{\hat{p}}$ and transcription index vector $\boldsymbol{I}$ are monotonic sequences ranging from 00 to $L-1$, a distance-aware attention reconstruction is performed by Eq. (9):
@@ II-C Attention reconstruction
where $d_{i,j}=(\hat{p}_{i}-I_{j})^{2},0\leq i\leq T-1,0\leq j\leq L-1$. $d_{i,j}$ represents the semantic distance between the $i$-th encoder frame and the $j$-th token. $\sigma$ is a learnable hyperparameter. It’s evident that as the distance $d_{i,j}$ decreases, the attention weight $\hat{\alpha}_{i,j}$ increases.
II-D Alignment predictor
As mentioned in Section II-A, an alignment predictor is required to generate alignments at inference stage. In EffectiveASR, the alignment predictor is built with two 1D-Convolution layers, each followed by layer normalization and ReLU function. During training, it takes acoustic embeddings $\boldsymbol{e}^{(\mathrm{s})}$ as inputs to generate a predicted alignment $\boldsymbol{\delta}^{*}$, while using the alignment $\boldsymbol{\delta}$ produced by the alignment generator as the label to update the predictor. In the inference stage, we use the alignment $\boldsymbol{\delta}^{*}$ produced by the alignment predictor to construct the attention matrix, as shown in the Fig. 1(b). At the same time, we accumulate the $\boldsymbol{\delta}^{*}$ and round it to predict the length of output tokens. The Mean Square Error (MSE) loss is used to learn the alignment, and the total loss is defined as:
$\mathcal{L}_{total}=\mathcal{L}_{\rm{CE}}(\boldsymbol{y},\boldsymbol{y^{*}})+\lambda\mathcal{L}_{\rm{MSE}}(\boldsymbol{\delta},\boldsymbol{\delta^{*}})\$ (10)
where $\lambda$ is the weight of alignment loss, and $\mathcal{L}_{\rm{CE}}$ is the cross-entropy (CE) loss between the text outputs and labels.
III Experiments
III-A Experimental Setup
The proposed method is evaluated on two public Mandarin corpus: 178 hours AISHELL-1 [22] and 1000 hours AISHELL-2 [23]. We use 80-channel filter banks computed from a 25 ms window with a stride of 10 ms as features. SpecAugment [31] and speed perturbation are used for data augmentation for all experiments. The sizes of the vocabulary for the AISHELL-1 and AISHELL-2 tasks are $4,233$ and $5,208$, respectively. The EffectiveASR model is built in base and large size. For both sizes, the layer numbers of {mel-encoder, text-encoder, alignment predictor, decoder} are set to {12, 1, 2, 6}, while the attention heads number of base model is $4$ with hidden dimension of $256$ and the attention heads number of large model is $6$ with hidden dimension of $384$. The hyperparameter $\lambda$ in Eq. (10) is set to $1.0$ and the learnable hyperparameter $\sigma$ in Eq. (9) is initialized with $0.5$. The proposed models are developed by ESPnet [32]  [^1] https://github.com/espnet/espnet.git, which is a popular open-source ASR toolkit. All experiments are performed on $8$ NVIDIA Tesla V100 GPUs, following the default configurations of ESPnet.